# Nuclear Fragmentation at the Future Electron-Ion Collider

### v2

@@ Abstract
We study low-energy nuclear physics at the future Electron-Ion Collider (EIC) at Brookhaven. By comparing the standard theory of electron-nucleus scattering with the equivalent photon method applied to Ultraperipheral Collisions (UPC) at the Large Hadron Collider (LHC) at CERN, we extract valuable insights into both processes. In the limit of extremely high beam energies and small energy transfers, very transparent equations emerge. We apply these equations to analyze nuclear fragmentation in UPCs at the LHC and $eA$ scattering at the EIC, demonstrating that the EIC could facilitate unique photonuclear physics studies.
@@
Figure 1: Cross sections for the production of fragments as a function of their masses in ultraperipheral collisions of 208Pb (open red squares) and 328U (closed blue circles) nuclei at the LHC.
@@
Figure 2: Cross sections for the production of fragments as a function of their masses for electron-208Pb (open red squares) and electron-328U (filled blue circles) collisions at the LHC.

### v3

@@ Abstract
We explore the potential of conducting low-energy nuclear physics studies, including nuclear structure and decay, at the future Electron-Ion Collider (EIC) at Brookhaven. By comparing the standard theory of electron-nucleus scattering with the equivalent photon method applied to Ultraperipheral Collisions (UPC) at the Large Hadron Collider (LHC) at CERN. In the limit of extremely high beam energies and small energy transfers, very transparent equations emerge. We apply these equations to analyze nuclear fragmentation in UPCs at the LHC and $eA$ scattering at the EIC, demonstrating that the EIC could facilitate unique photonuclear physics studies. However, we have also shown that the fragmentation cross-sections at the EIC are about 1,000 times smaller than those at the LHC. At the LHC, the fragmentation of uranium nuclei displays characteristic double-hump mass distributions from fission events, while at the EIC, fragmentation is dominated by neutron emission and fewer few fission products, about 10,000 smaller number of events.
pacs:
@@
Figure 1: Cross sections for the production of fragments as a function of their masses in ultraperipheral collisions of 208Pb (open red squares) and 238U (closed blue circles) nuclei at the LHC.
@@
Figure 2: Cross sections for the production of fragments as a function of their masses for electron-208Pb (open red squares) and electron-238U (filled blue circles) collisions at the LHC.
@@
At the EIC, the integrated luminosities for electron-nucleus collisions are expected to range between $10^{33}-10^{34}$ cm2s-1. Assuming a luminosity of ${\cal L}=10^{34}$ cm2s-1, approximately $10^{8}$ neutrons per second can be produced from electron-uranium nucleus fragmentation. Event rates for the production of uranium fission fragments are anticipated to fall within a similar range. These neutrons and fragments will predominantly travel in the forward direction with negligible transverse momenta, making their detection challenging without specialized equipment. A possible solution is to utilize a setup similar to the Zero Degree Calorimeter (ZDC) employed with the ALICE detector at CERN ARNALDI2006235 . The specifics of detection techniques, however, lie beyond the scope of this work goto19 . Notably, many low-energy processes in relativistic heavy-ion collisions were predicted as early as the 1980s BERTULANI1988299 when suitable detectors were unavailable, yet they were eventually observed experimentally decades later. It is highly plausible that the low-energy processes described in this study will also become experimentally accessible at the EIC.